OpenStudy (anonymous):

the lengths of two diagonals of a particular rhombus are 10 ft and 14 ft, what is the perimter of the rhombus?

OpenStudy (anonymous):

since diagonals of a rhombus bisect each other at right angles, so thie halves form a right angled triangle with the side being the hypotenuse

OpenStudy (anonymous):

so side^2 = 7^2 + 5^2 side^2 = 49 + 25 = 74 so side = sqroot(74) = 8.6 ft approx

OpenStudy (anonymous):

That's what I was getting, but I wasnt sure if 8.6 was the right answer. Didn't know if it had to be a whole number or not.

OpenStudy (anonymous):

do you multiply 8.6 times 4 because it's a rhombus?

OpenStudy (anonymous):

yes perimeter will be = 4*side if u want more exact answer, then perimeter = 4*sqroot(74) = \[4\sqrt{74}\]

OpenStudy (anonymous):

Thank you!